### v1

# A Complete Active Space Self-Consistent Field (CASSCF) approach for molecules in QED environments

Abstract
Multireference systems are usually challenging to investigate using ab initio methods as they require an accurate description of static electron correlation. The urgency of developing similar approaches is even more pressing when molecules strongly interact with light in quantum-electrodynamics (QED) environments. In fact, in this context, multireference effects might be induced or reduced by the presence of the field. In this work, we extend the Complete Active Space Self-Consistent Field (CASSCF) approach to polaritonic systems. The method is tested on benchmark multireference problems and applied to investigate field-induced effects on the electronic structure of well-known multiconfigurational processes. We analyze the strengths and limitations of the method with particular attention to the appearance of possible origin dependencies.
Perugia] Dipartimento di Chimica, Biologia e Biotecnologie, Università degli Studi di Perugia, Via Elce di Sotto, 8,06123, Perugia, Italy Trondheim] Department of Chemistry, Norwegian University of Science and Technology, 7491 Trondheim, Norway Perugia] Dipartimento di Chimica, Biologia e Biotecnologie, Università degli Studi di Perugia, Via Elce di Sotto, 8,06123, Perugia, Italy
1 Introduction
Manipulation of molecular properties by quantum fields is becoming nowadays a very promising non-intrusive alternative to chemical controls 1, 2, 3, 4, 5, 6, 7, 8, 9. The observation of these effects was possible thanks to the significant technological advances made in the fabrication of efficient optical devices (optical 10, 11, plasmonic 12, 13 and superconducting 14, 15, 16 cavities, waveguides 17, 18, 19, etc.). The simplest of these devices is the Fabry-Pèrot optical cavity, device composed of two mirrors whose separation, L, is related to the wavelength of the field inside the cavity. In these conditions, energy is exchanged between the system and the cavity field, inducing a phenomenon called quantum Rabi oscillations. Due to the strong light-matter interaction, photons and molecular states lose their individuality and show mixed molecular and photonic character, generating hybrid states known as ”polaritons” 20, 21, 22, 23. The formation of polaritons is associated with a splitting of the energy levels, known as Rabi splitting (), that increases with the strength of the light-matter coupling () (see Fig. 1).
It is noteworthy to point out that molecules can interact strongly with the field also in the absence of real photons, through a direct coupling between the molecular states and the vacuum fluctuations of the electromagnetic field. In the past decade, a lot of research was devoted to explore the potential of strong light-matter interaction on a plethora of physical or chemical phenomena. In particular, by coupling molecules or materials with different optical devices it was possible to manipulate the absorption, emission, and scattering properties of molecules and materials 24, 25, 26, 27, to induce Bose–Einstein condensation of polaritons 28, 29, 30, to increase conductivity in organic semiconductors 31, or to manipulate the spin properties of matter 32, 33, 34. One of the most debated but also interesting aspects was observed in chemistry, where field-induced effects on chemical reactivity were demonstrated to inhibit, catalyze, and even make some reactions selective toward specific products 35, 36, 37, 38.
Despite the large attention devoted to this field in the past decade, the physics behind many of these processes still remains unclear, and progresses are still very slow due to the many limitations in the validation of the experimental setup 39, 40, 41. For these reasons, the development of effective theoretical approaches is crucial to overcome experimental limitations and to propose a consistent description of these phenomena 42, 43. In recent years, a small number of research groups started developing some extensions to the most commonly used ab initio methodologies, i.e., Density Functional Theory (DFT) 44, 45, 46, Hartree-Fock (HF), Full Configuration Interaction (FCI), Coupled Cluster (CC) 47, 48, 49, 50, 51, and second-order Møller-Plesset perturbation theory (MP2) 52, to investigate coupled electron-photon systems. Recently, fully relativistic versions of QED-HF and QEDFT have been also proposed to investigate molecular systems containing heavy atoms in strong coupling conditions 53, 54. The quantum electrodynamics (QED) extensions to HF and DFT are quite efficient and can be easily applied to relatively large systems. However, they completely miss correlation or describe it in a very approximated way. Moreover, in the DFT case, an effective exchange-correlation functional for the electron-photon interaction is still far from being formulated 42.
Coupled Cluster theory is known to be the most reliable method in quantum chemistry to simulate molecular systems and also for interacting electron-photon systems, its QED extension is at the moment the reference ab-initio methodology. However, as for the standard electronic theory, the applicability of this method is limited to small to medium sized molecules and its description becomes qualitatively poor when multiconfigurational effects come into play. This problem is well-known and it prevents the investigation of a relatively large class of complexes, including several systems involved in chemically and biologically interesting precesses 55. Belong to these systems, open-shell transition metal complexes, (poly)radicalic systems, excited states of organic molecules, and torsion or breakings of conjugated bonds 56, 57.
The state-of-the-art reference approaches usually used to investigate these systems are the Multiconfigurational Self-Consistent Field (MCSCF) methods, which are based on a wavefunction defined as linear combination of a reduced number of determinants or Configuration State Functions (CSFs). Nowadays a wide variety of MCSCF methods have been developed. The main difference between them lies in the protocol employed to select the determinants to include in the expansion 58. The most popular MCSCF method is the Complete Active Space Self-Consistent Field (CASSCF), which is based on a partitioning of the orbital space into three classes: inactive, virtual, and active orbitals. The multideterminantal expansion in the case of CASSCF theory is obtained by choosing a certain number of so-called active electrons and computing all the possible excitations of these electrons within the active space. If the active space is chosen properly, (i.e., the most correlated orbitals and electrons are included)the CASSCF wavefunction is able to provide a qualitatively-correct description of the molecular system. Despite this method is equivalent to a FCI expansion within the active space, with the current computational facilities, it can be applied, in a relatively efficient way, to active spaces with up to 20 electrons in 20 orbitals 59.
In the latest years, some of the most common ab-initio multireference methods have been extended to coupled-light matter systems 60, 61, 62. In particular, the implementation of QED-CASCI and QED-DMRG provided the first steps towards the inclusion of static correlation effects in optical cavities. However, the CASCI procedure does not optimize the orbitals, leading to variationally higher energies with respect to its SCF counterpart 63, and although the DMRG method has a polynomial scaling, its cost is higher than CASSCF for small active spaces.
In this work, we present the first extension of CASSCF theory to coupled light-matter systems (QED-CASSCF). The implementation is performed following a restricted-step second order optimization scheme, according to the implementations in Refs. 64, 65, 66, 67.
This paper is organized as follows: in Section 2, a brief summary of the CASSCF method and its second-order implementation is presented. In Section 3, the fundamental details of QED-CASSCF are described. In Section 4, QED-CASSCF is tested on small molecules against other ab-initio methodologies. In this section, QED-CASSCF is also used to investigate the field effects on well known multireference chemical processes. We conclude the work with future perspectives presented in Section 5.
2 CASSCF Theory
In MCSCF theory the wavefunction is written as the linear combination of a certain number of Slater determinants or Configuration State Functions (CSFs)
and it is optimized by variationally minimizing the energy with respect to the molecular orbital (MO) and configuration interaction (CI) coefficients ( and , respectively)
Among the MCSCF methods 68, 69, 70, 71, 72, the Complete Active Space Self-Consistent Field (CASSCF) approach is still the state-of-the art procedure for molecular systems dominated by static correlation. The orbital space of a molecule in the CASSCF approximation is partitioned into three subspaces, as depicted in Fig. 2: inactive orbitals are always fully occupied, active orbitals have non-integer occupations, and virtual orbitals are always empty. In this work the different orbitals are addressed according to the following indices choice: for inactive orbitals, for active orbitals, for virtual orbitals, and for generic orbitals.
The multideterminantal expansion for the wavefunction is obtained considering all the determinants computed by distributing active electrons in active orbitals, namely CAS(N,M).
The first implementations of CASSCF relied on first-order solvers and among these, the super-CI approach is the most widely applied 73, 74. However, using this scheme, the optimization of the CASSCF wavefunction is affected by severe convergence issues. This led to the development of robust second-order quadratically-convergent solvers which exhibit better convergence properties in a reduced number of iterations 64, 75, 76, 77, 78, 79, 80, 81, 82, 83, 67. CASSCF implementations based on second-order algorithms are nowadays relatively well-established in the quantum chemistry community and are available in the majority of the mostly used software packages 84, 85, 86, 87, 88. This approach is particularly convenient for the application to simulate small molecular systems, and calculations on larger systems are nowadays efficiently performed with the aid of parallelization, AO-based implementations, or by exploiting the Cholesky decomposition or the resolution of identity (RI) approximation of the two-electron integrals 83, 67.
In MCSCF theory, the first step towards the implementation of the method is an appropriate choice of the parametrization for the wavefunction. The orbital part is parametrized in terms of an exponential transformation with the anti-Hermitian operator defined as:
where is the spin traced singlet excitation operator. While no alternatives to the parametrization of the orbital space are available, the CI space is liable to multiple choices depending on the CASSCF variant that has to be implemented. In the case of State-Specific (SS) CASSCF method, a linear parametrization for the CI part is commonly adopted and the wavefunction can be written as 64:
where is indicated as the current expansion point (CEP) and constitutes the current approximation to the wavefunction. , instead, is a projector operator that removes the components parallel to the CEP from .
The choice of the exponential parametrization allows to consider only the non-redundant rotations between the orbital spaces (i.e., orbital rotations which contribute to the energy: inactive-active, inactive-virtual, and active-virtual). The linear parametrization of the CI space contains only one redundant parameter when is parallel to the reference state . However, this is not an issue as this redundancy can be easily tracked and it does not interfere with the optimization procedure.
In the following of this section we will quickly recap the implementation details of the second-order SS-CASSCF algorithm that will be applied, in the next section, for the extension to strongly coupled electron-photon systems.
2.1 Second-Order CASSCF
The implementation of a second-order algorithm is based on the definition of a quadratic model to locally expand the energy up to the second order in :
where indicates a small variation of the wavefunction parameters (, ). In Eq. 5, is the CASSCF energy at the current expansion point. is the gradient vector containing the first derivatives of the energy with respect to the configuration (c) and orbital (o) coefficients calculated at the current expansion point:
Finally, is the Hessian matrix containing the pure configuration (cc), orbital (oo), and the mixed configuration-orbital (co) and orbital-configuration (oc) parts:
Also the Hessian matrix elements are calculated at the current expansion point.
In principle, a second order algorithm requires the explicit computation of the electronic Hessian. However, this procedure can be performed efficiently in a direct fashion. This approach passes through the calculation of the so-called sigma vectors defined as:
Here, are the so-called trial vectors used in the Davidson scheme 89 which contain the wavefunction parameters with respect to the orbital and configuration parts
This approach allows for large scale operations without explicitly calculating and storing the blocks of in memory 90. Moreover, sigma vectors can be conveniently computed applying minor modifications to intermediate quantities already calculated for the gradients.
By minimizing the quadratic function defined in Eq. 5 the equation for the Newton step is obtained as:
This procedure is not optimal since it requires the Hessian to be positive-definite. This condition is not always guaranteed, especially at the beginning of the optimization procedure. Even when it is respected, the computed step may not point in the direction of the minimum, leading to a poor convergence 91. This problem can be largely reduced by adopting a restricted step procedure 64, 92. According to this scheme, the choice of the step is restricted within a well-defined trust region with radius, :
where we assume the potential energy surface to be correctly approximated by the quadratic model. This ensures the Hessian to be positive-definite. This procedure was first described by Levenberg and Marquardt 93, 94 and is based on the solution of the level-shifted Newton equation:
where the level-shifting parameter corresponds to the lowest eigenvalue of the augmented Hessian:
The relation between the eigenvalues of and is given by the Hylleraas-Undheim-MacDonald theorem 95, 96:
Therefore, the choice of the lowest eigenvalue of the augmented Hessian as a level-shifting parameter guarantees to be positive-definite. This approach requires, in principle, to solve the linear system in Eq. 12 at every iteration. However, the inversion of could be computationally inefficient. An alternative way can be obtained by diagonalizing the augmented Hessian in Eq. 13. This can be done efficiently in an iterative fashion by implementing the sigma vectors in Eq. 8. Once the lowest eigenvalue and eigenvector are obtained, the step for the new iteration can be computed as:
where is the eigenvalue of the Hessian and is a scaling parameter that forces the step to lie withing the trust region. The trust radius is changed adaptively during the procedure following the strategy described by Fletcher in Ref. 91. As long as the step is kept within the trust region, the Newton’s step in 10 is recovered and convergence to the closest minimum is guaranteed.
The presented trust-region optimization method can be summarized as follows:
1.
Calculation of the gradient by Eq. 6. If its norm is lower than a certain threshold, convergence is reached;
2.
Check the step and adjust the trust radius according to the procedure described by Fletcher in Ref. 91;
3.
Iteratively diagonalize the augmented Hessian in Eq. 13;
4.
Update the wavefunction parameters (Eq. 15) and return to step 1.
3 QED-CASSCF
The description of molecular systems in optical cavities requires to account explicitly for the quantum character of the electromagnetic field. In our work we will use the well-established non-relativistic Pauli-Fierz Hamiltonian in length gauge and dipole approximation as starting point for the development of the QED-CASSCF approach 47:
Here is the dipole moment operator, is its expectation value and is the light-matter coupling parameter that depends on the quantization volume and the polarization vector . For simplicity a single cavity mode is included in our treatment, but extension to more modes can be applied in a trivial manner. To ensure origin invariance, Hamiltonian 16 has been already expressed in the coherent-state basis 47. Following the notation used in Ref. 98, we can conveniently set and rewrite the Pauli-Fierz Hamiltonian as
In this form the dipole-self energy term is included into the one- and two-electron integrals
with and being the electric dipole and quadrupole integrals, respectively, and is the number of electrons in the molecule.
For a general QED-MCSCF theory the wavefunction can be expressed as 60:
where is the photonic state of the cavity mode. In this case the CI coefficients refer to the non-interacting electron-photon states defined as
In this way a generic polaritonic state can be written as
with representing the ground state.
For the sake of readability of the following section we conveniently separate the Pauli-Fierz Hamiltonian into three terms:
where we need to remember that contains the one- and two-electron integrals modified for the dipole self-energy term. From now on, we will refer to the Pauli-Fierz Hamiltonian as unless explicitly stated.
The extension of CASSCF to molecular polaritons and its second-order implementation requires the computation of modified energy contributions, gradients and sigma vectors. The expression for the QED-CASSCF energy can be written as:
where
has the same form of the purely electronic CASSCF energy 83 with
being the one- and two-body reduced density matrices, respectively, and
the inactive Fock matrix. The purely photonic contribution to the energy is given by:
where is the number of photons in the photonic state of the cavity mode. Finally, the contribution by the bilinear coupling is written as:
The general expression for the CI gradient is given by:
where is the QED-CASSCF energy at the current expansion point. Similarly to the energy, also the CI gradient can be split up into different terms:
where
The bilinear contribution to the CI gradient:
is made of two ”off-diagonal” contributions
and
The general expression for the orbital part of the gradient is given by:
The purely electronic contribution is known 58:
where
is the generalized Fock matrix. It is easily verified that the purely photonic contribution to the orbital gradient is zero since and commute. Finally, the interaction term to the gradient reads as:
where the one-body density matrices are defined as:
Finally, the sigma vectors defined in Eq. 8 can be calculated as:
where is the one-index transformed Hamiltonian defined as 58:
where
4 Results
In this section, QED-CASSCF will be tested on small multireference systems and used to investigate field-induced multiconfigurational effects. All the calculations are performed with a development version of the eT software package 97.
4.1 Dissociation of the N2 ground state
The homolytic dissociation curve of N2 is a typical test case for investigating the behavior of multireference methods. For this system, the active space contains the six valence electrons inside the six p orbitals, i.e., CAS(6,6). The calculations were performed using a Dunning’s cc-pVDZ basis 99, light-matter coupling a.u., cavity frequency eV, and the field polarization along the axis, coincident with the main molecular axis. All the QED calculations are performed including a single photonic state. The potential energy curve has been computed over 32 steps from to Å. In Fig. 3a, the potential energy curve computed at the QED-CASSCF level is compared against QED-CASCI and QED-CCSD.
For distances up to 5Å, the energy profiles are in qualitative agreement with the purely electronic potential energy curves (PECs), with coupled cluster exhibiting lower energies with respect to CAS-type energies. Notice, however, that, at Å, CC, after showing an unphysical maximum around Å, stops converging due to the emergence of multireference effects associated with the homolytic dissociation of the system. Despite showing an higher energy for the dissociation plateau compared to both QED-CC and QED-CASSCF, QED-CASCI describes the right energy profile in a wide range of energies as expected already for the bare electronic system. The higher energy of the dissociated system can be clearly explained by the lack of dynamical correlation (included in QED-CC) and by the use of unoptimized molecular orbitals (present in QED-CASSCF). However, it is important to highlight the presence of an unphysical jump in energy between and Å which in not registered for the purely electronic profile (see Fig. S1 in Supporting Information). This behavior is indeed given by the guess used for the calculation. In fact, while the purely electronic Hartree-Fock potential energy curve is continuous in the whole energy range, the QED-HF curve shows a drastic energy change at the same internuclear distances (see Fig. 3b). This is also associated to a breaking of the orbitals degeneracy. This problem is fixed by QED-CASSCF, as shown in Fig. 3. In this case, indeed, the variational optimization of the orbitals avoids discontinuities due to the poor quality of the guess, predicting a continuous energy profile in the whole distances range. This behavior is a demonstration that the presence of the field can, in some cases, complicate the simulation of molecular systems characterized by multireference effects.
In order to estimate the entity of the field-induced effects, we also computed the energy difference between the different QED methods and the corresponding electronic ones (see Fig. 4). As expected, the energy difference goes to a plateau for CASSCF (see Fig. 4a) while the same discontinuity is observed for CASCI and HF (see Fig. 4b) . Such difference is particularly large for the Hartree-Fock case where it amounts to for the chosen coupling strength. This energy contribution is entirely given by the dipole self-energy term, that it the only contribution affecting the QED-HF energy 47. These results clearly highlight the importance of orbital optimization to obtain qualitatively-correct results especially in the cases where multireference effects become dominant.
Before concluding, it is important to stress that even for very large coupling values (per molecule) like those applied in this test, the field effects on the ground state energy are very small. Much more significant effects could be observed on the excited states energies, but their study goes beyond the scope of this paper and will be analyzed in details later in a follow-up work.
4.2 Field-induced effects on the torsion of ethylene
The description of bond-breakings, torsions along conjugated bonds, and conical intersections has always been a challenge for single-reference methods 100. CASSCF is still the most-commonly employed method to provide a good guess for the description of such phenomena. In this section we investigate the double bond torsion of an ethylene molecule coupled to an optical cavity. This molecular system has been largely studied in the literature 101. This work does not intend to provide any rigorous quantitative treatment of the chemistry of the system but it only shows the effects induced by the cavity environment.
The torsion of the double bond induces interesting properties to the electronic structure of ethylene in particular when the dihedral angle approaches 90°. For this particular geometry, the two partially occupied (frontier) orbitals are practically degenerate and have occupation numbers equal to 1. This configuration confers an high multiconfigurational character to the system. In this section we focus our attention on the effects induced by the field on the frontier orbitals energies for ethylene in the ° configuration. The calculations were performed using a cc-pVDZ basis set and a minimal active space containing the two electrons in the and orbitals, CAS(2,2). In this case, the cavity frequency has been chosen to match the excitation energy to the lowest bright electronic transition of the molecule ( eV). The polarization direction has been directed along the main molecular axis (). The orbital energy profiles were recorded by increasing the light-matter coupling from to a.u. (see Fig. 6). When the coupling is increased the orbital energies remains almost constant until coupling values of about 0.25 a.u. to then change rapidly and reaching a gap of eV (see Fig. 6a). This breaking of the orbital degeneracy, with consequent reduction of the multiconfigurational character of the system, reflects also on the orbitals’ occupations (see Fig. 6b). Now, one of the frontier orbitals appears occupied with more than one electron while the other is occupied with small fractions of an electron. This behavior is a consequence of the field-induced mixing between the two orbitals (as can be seen from Fig. 6c).
Summarizing, in this paper we observed that in certain conditions the field may induce changes on the multiconfigurational character of a molecular system. The analysis, performed in this work, has been conducted with light-matter coupling values that are way larger than those currently obtainable in actual experiments. However, if systems with smaller excitation energies are investigated and coupling the method with strategies to properly account for collective effects, our methodology have the potential to become a reference approach to simulate field induced effects on the electronic structure of multireference system.
4.3 QED-CASSCF origin dependence
It is now well established (as demonstrated in Ref. 47) that QED-HF theory has an intrinsic origin dependence if charged systems are investigated. This origin dependence does not appear on the total energy, because of the origin invariance of Hamiltonian 16 but it is very evident on the molecular orbitals and on their energies. This behavior has been addressed in detail by Riso et al. in Ref. 102. In particular they developed an alternative methodology know as strong coupling quantum electrodynamics Hartree Fock (SC-QED-HF) 102, 103 with the intent of constructing a consistent molecular orbital theory for molecules in strong coupling conditions. In this framework, the wavefunction is transformed by a unitary transformation and the orbitals are defined within an electron-photon correlation basis obtained by diagonalizing the operator . This approach provides fully origin-invariant molecular orbitals and is able to recover part of electron-photon correlation energy. Therefore, SC-QED-HF is better fit for the investigation of cavity-induced effects and provides a physically meaningful guess for performing post-HF calculations on charged systems. Recently, the strong coupling formalism has also been extended to second-order Møller-Plesset theory and response theory 52, 104.
Because of an explicit dependence of the Fock matrices on the dipole matrix elements that reflects also on the orbital gradient (as evident from Eq. 39) we expect also the QED-CASSCF molecular orbitals to be origin dependent for charged systems. It is important to stress that, in this case, since CASSCF explicitly depends on the orbitals, also the total energy will not be origin-invariant. Moreover, since a re-optimization of the orbitals is performed, we expect to observe this behavior also if SC-QED-HF is used as a reference. In this section we address this point in details. To prove this, we performed a simple test case on the hydroxide ion. Here, CAS(2,2) calculations were performed while shifting the molecule along the axis (the origin of the reference system has been centered on the oxygen atom). The cavity frequency was set, in this case, to eV and the light-matter coupling to a.u. Fig. 7 displays the expected variation of the total energy as a function of the shift along the axis. We demonstrated also, as stated earlier, that the computed energies remain unmodified regardless of the starting guess (SC-QED-HF or QED-HF).
5 Conclusions
In this work we have presented the first (at least to our knowledge) extension of the complete active space self-consistent field method to account for the effect of quantized electromagnetic fields in optical cavities, (QED-CASSCF). The implementation of the method was performed according to a restricted-step second-order optimization method. This strategy allows to overcome some of the convergence issues that characterize the CASSCF wavefunction optimization 83, 67.
The developed method was tested on small multireference molecules where the inclusion of static correlation is crucial to provide at least a qualitative description of the system. The comparison between QED-CASSCF and QED-CASCI for the dissociation curve of the nitrogen molecule proved the importance of the orbital optimization to avoid unphysical behaviors arising from possible poor quality guesses.
We also investigated the effects induced by the cavity field on the ethylene torsion process. Our calculations show that the presence of the electromagnetic field can remove the degeneracy between the and orbitals reducing, in specific coupling conditions, the multireference character of the system.
Finally, we addressed the origin dependence of our method in case of charges systems. In particular, the simple case study of the OH- ion showed that, even if an origin-invariant reference is used, the method still preserve a dependence on the choice of the reference system. This issue will be solved in a future paper where the Strong Coupling transformation proposed in Ref. 102 will be applied to remove the origin-dependence of QED-CASSCF also for charged systems.
It is important to remind that, the results obtained in this paper, due to the very large coupling values we used, are not directly comparable with experimental results. However, the aim of the paper, was not much to provide accurate simulations of molecular systems in optical cavities, but more proposing a new methodology able to investigate multireference systems coupled to electromagnetic fields. In a near future this methodology will be coupled with strategies to include collective effects that will allow for more realistic simulations 105, 106, 107.
With this work we move one step further towards an accurate description of multiconfigurational molecular systems in optical cavities. This paves the way for the investigation of chemically/photochemically-relevant processes that until this day have been investigated through model Hamiltonians or other approximated methods. {acknowledgement} R.A and E.R. acknowledge funding from the European Research Council (ERC) under the European Union’s Horizon Europe Research and Innovation Programme (Grant n. ERC-StG-2021-101040197 - QED-SPIN). H.K. acknowledges funding from the European Research Council (ERC) under the European Union’s Horizon 2020 Research and Innovation Programme (grant agreement No. 101020016). {suppinfo}
The SI contains plots with direct comparison between potential energy surfaces calculate with QED and bare electronic methods.
5.1 Code availability
The program used to perform the calculations shown in this work is available from the corresponding author upon reasonable request.